A ball is thrown vertically upward attains max height of 45 m.Thr time after which velocity of the ball becomes equal to half the velocity of projection is.(g =10m/s2)

Grade:

1 Answers

Mayank Ranka
2 years ago
Let the velocity of projection be 'u'. Then 'v' = 0 (at top height)


By equation of motion, v²=u²+2as
i.e.,
v2= u2- 2gh (Since here s = h and when ball is thrown accealaration is taken as -g, i.e., a=-g)

u =2gh−−−√

u = 30 m/s

now, time, when v'= 15 m/s

from first equation of motion: v= u+at
Here a=-g
So,
v = u -gt
15 = 30 - 10t

10t = 15

t = 1.5 s

So after 1.5 seconds velocity becomes hal